OpenStudy (anonymous):

Simplify and solve this equation for q: 3q + 5 + 2q – 5 = 65

OpenStudy (anonymous):

the fives cancel out and you are left with 5q=65 Dividing by 5, you get, q=13

OpenStudy (haseeb96):

3q +2q -5 +5 -65=0 first add the same terms 5q -65=0 5q =65 q=65/5 q=13 is the correct answer